# On the long-time limit of the mean curvature flow in closed manifolds

### v1

@@ 2.2. Ilmanen’s regularity theorem
Although it is the case that the flow preserves the topology of $M$ in this sense it could be the case that there is strict genus drop in the limit as $t\to\infty$. In the following we focus only on the connected component $\tilde{M_{t}}$ of $M_{t}$ homeomorphic to $M$, and denote its corresponding limiting stable minimal surface $\Sigma$; the reason this suffices for us because, by the $\pi_{2}$ triviality assumption and that the other components pinched off the flow will be spheres, at any given time $M_{t}$ will be homotopic to just this component. Now, if the convergence is with multiplicity one then it will be the case $\Sigma$ is orientable and isotopic to $M$. Supposing this is not the case, then from the proof of Theorem 1.4 we have that either $\Sigma$ is nonorientable and $M_{t}$ converges to $\Sigma$ with multiplicity 2, or $M_{t}\to\Sigma$ with some multiplicity $m$ and there is at least one bridge point $p$, within which for some large time $s$ and some small $r>0$ we have $M_{s}\cap B(p,r)$ has less than $m$ components – our task is to rule out this latter case. Taking $r$ slightly larger if necessary $M\cap(B(p,2r)\setminus B(p,r))$ is comprised of $m$ disjoint annuli, where $m$ is the multiplicity of convergence; after a slight isotopy we may suppose these annuli $A_{i}$ lay in parallel planes and so may order them from "top" to "bottom" in the obvious sense, with $A_{1}$ being the top annulus. Without loss of generality the top annulus at the point $p$ is nontrivial in the sense it doesn’t border a disc in $M_{s}\cap B(p,r)$; if not below we can consider the next lowest disc and so on.
@@
Considering one of the boundary components $\gamma$ of $A_{1}$, we may clearly find an embedded disc $D$ in $N$ disjoint from $M_{s}$ with boundary $\gamma$. If $\gamma$ is nonseparating or splits $M_{s}$ into two components of genus greater than zero then $D$ will be a compressing disc because $\gamma$ will be homotopically nontrivial in $M$, which gives a contradiction. If the number of bridge points connected to the top sheet (i.e. where $B(p_{i},r_{i})$ are the bridge points and neighborhoods, the connected component of $M\setminus\cup_{i}B(p_{i},r_{i})$ containing $A_{1}$) is greater than one then $\gamma$ is nonseparating, so that we are done; let us suppose that $p$ is the only bridge point where the top sheet connects to the lower sheets then. If $\Sigma$ is not an embedded sphere or projective plane then $\gamma$ would disconnect $M$ into two components of nonzero genus, completing the argument in this case. If $\Sigma$ is a sphere or projective plane a bit more arguing is required. As before first suppose that $\Sigma$ is a sphere. Then because $N$ has trivial $\pi_{2}$, we may homotope $\Sigma$ and hence $M_{s}$ to lay in a small coordinate chart; of course because ${\mathbb{R}}^{3}$ is contractible $M_{s}$ in this case cannot be incompressible. If it is a projective plane again the boundary of its tubular neighborhood is a sphere, and we may proceed as in the first case. ∎
@@
Leaving the world of stable minimal surfaces we now consider the use of our first convergence statement, Theorem 1.1, to essentially replace the tightening process in minmax. Recalling that in the setting of Corollary 1.10 setting $N\simeq S^{3}$ we may consider the standard one parameter sweepout $\Sigma_{s}$ of it by $2$–spheres, where there are two singular values of the sweepout that are single points corresponding to the north and south poles under a fixed choice of diffeomorphism from $N$ to $S^{3}$. Perhaps the most natural thing to do then is to consider the flow of the sweepout under the flow, by which we mean the family of the flows $(\Sigma_{s})_{t}$ of the individual leaves $\Sigma_{s}$. Recalling that the width of any sweepout of $S^{3}$ by spheres is positive by the isoperimetric inequality, we see if we could show that if the flow will remain a sweepout, then Theorem 1.1 says that the flow of at least one leaf will not go extinct, and because $S^{3}$ is simply connected and embedded hypersurfaces in such spaces are orientable Proposition 3.7 gives the limit will be an embedded minimal sphere, as usual potentially with multiplicity. There are some issues with the approach though, unfortunately.
@@
The biggest issue in general appears to be that the level set flow of some of the leaves may fatten, but in our particular setting this may be ruled out as we discuss below. There is also an issue with preserving the typical regularity assumptions one make in the definition of sweepout. Recall that a sweepout of $N$ (here, $1$–parameter) is a family of surfaces $\Sigma_{s}$, $s\in[0,1]$, which for finitely many $s$ is singular in a finite set of points and smooth elsewhere and whose union is all of $N$. Now, we know that for a fixed leaf $\Sigma_{s}$ the flow $(\Sigma_{s})_{t}$ will be smooth for almost all times $t$, but because we wish to consider the flow of the family, parameterized over the whole unit interval (which is uncountable so Baire category theorem cannot apply), it doesn’t seem clear if we can show the flow of the whole family will be smooth for almost all or even many times. The current state of the art also doesn’t rule out the size of the singular set of a flow being greater than a finite union of points, either.
@@
With respect to the discussion above one may be able to proceed by generalizing the results of minmax theory to more general sweepouts including those "generated" by the flow but instead we will take a somewhat more hands on approach. Denote by $E(s)$ the extinction time of the flow of $\Sigma_{s}$; if this never occurs of course we set $E(s)=\infty$ so it is a function into the extended reals. We next claim that $E(s)$ is sequentially continuous. To see this, we first note that because each of the $\Sigma_{s}$ are spheres their singularities under the flow must be mean convex and so each of their flows is nonfattening – implying for instance that the almost regular flows $(\Sigma_{s_{i}})_{t}$ are well defined. Fixing an $s_{0}\in[0,1]$, nonfattening also implies that for a sequence $s_{i}\to s_{0}$, after potentially passing to a subsequence by Brakke compactness, $\lim\limits_{s_{j}\to s_{0}}(\Sigma_{s_{i}})_{t}\to(\Sigma_{s_{0}})_{t}$ as Brakke flows because $\lim\limits_{s_{j}\to s_{0}}(\Sigma_{s_{i}})_{t}$ is a weak flow associated to $\Sigma_{s_{0}}$. With this in mind suppose then that there was a sequence $s_{i}\to s_{0}$ and $\epsilon>0$ for which $E(s_{0})<A-\epsilon$, where $A=\lim\limits_{i\to\infty}E(s_{i})$.
@@
At time $T=E(s_{0})$, by Colding and Minicozzi [13] $(\Sigma_{s_{0}})_{T}$ is the union of a finite number of compact embedded Lipschitz curves corresponding to cylindrical singularities along with a countable collection of points corresponding to spherical singularities. This gives that for $i$ sufficiently large we make take the mass of $(\Sigma_{s_{i}})_{T}$ to be as small as we like, so in particular the clearing out lemma says these flows must go extinct before $t=A-\epsilon/2$, giving a contradiction to show that $E(s)$ is upper semicontinuous. Similarly $E(s)$ is lower semicontinuous, giving our claim of sequential continuity.
@@
Because $E(s)$ is sequentially continuous, by the compactness of the unit interval $\sup\limits_{s\in[0,1]}E(s)=T$ is obtained for some time $s_{0}$, recycling notation. Now as above we have $(\Sigma_{s_{0}})_{T}$ must consist of the union of a finite number of compact embedded Lipschitz curves corresponding to cylindrical singularities along with a countable collection of points corresponding to spherical singularities. Picking some smooth time $t^{*}$ right before $T$, the resolution of the mean convex neighborhood conjecture gives that $(\Sigma_{s_{0}})_{t_{*}}$ is mean convex and after rescaling is locally modeled on bowl solitons, round cylinders, or round spheres.
@@
To see why this is useful, first note that $s_{0}\in(0,1)$: this is because for $s$ very near 00 or $1$ that $\Sigma_{s}$ are approximately small concentric round spheres contained in a coordinate chart so that $\lim\limits_{s\to 0,1}E(s)=0$; on the other hand for a fixed smooth $\Sigma_{s}$ its smooth flow exists for a short but strictly positive time. Secondly, the flows $(\Sigma_{s})_{t}$ are pairwise disjoint in $s$ because $\Sigma_{s}$ are and almost regular flows are weak set flows. In particular, for $s-s_{0}$ sufficiently small $s$ Brakke regularity says that $(\Sigma_{s})_{t^{*}}$ is mean convex and a graph over $(\Sigma_{s_{0}})_{t_{*}}$ laying on one side or the other of it, and we can find $s$ so that the flow at time $t^{*}$ is on either side. Hence we see that for $s$ either slightly greater or smaller than $s_{0}$ $(\Sigma_{s})_{T}$ must not be singular, or in other words that their flow lasts at least slightly after time $T$ giving a contradiction to the definition of $T$. ∎

### v2

@@ 2.2. Ilmanen’s regularity theorem
Although it is the case that the flow preserves the topology of $M$ in this sense, it could be the case that there is strict genus drop in the limit as $t\to\infty$. In the following we focus only on the connected component $\tilde{M_{t}}$ of $M_{t}$ homeomorphic to $M$, and denote its corresponding limiting stable minimal surface $\Sigma$; the reason this suffices for us because, by the $\pi_{2}$ triviality assumption and that the other components pinched off the flow will be spheres, at any given time $M_{t}$ will be homotopic to just this component. Now, if the convergence is with multiplicity one, then it will be the case that $\Sigma$ is orientable and isotopic to $M$. Supposing this is not the case, then from the proof of Theorem 1.4 we have that either $\Sigma$ is nonorientable and $M_{t}$ converges to $\Sigma$ with multiplicity $2$, or $M_{t}\to\Sigma$ with some multiplicity $m$ and there is at least one bridge point $p$, within which for some large time $s$ and some small $r>0$ we have $M_{s}\cap B(p,r)$ has less than $m$ components – our task is to rule out this latter case. Taking $r$ slightly larger if necessary $M\cap(B(p,2r)\setminus B(p,r))$ is comprised of $m$ disjoint annuli, where $m$ is the multiplicity of convergence; after a slight isotopy we may suppose these annuli $A_{i}$ lay in parallel planes and so may order them from "top" to "bottom" in the obvious sense, with $A_{1}$ being the top annulus. Without loss of generality, the top annulus at the point $p$ is nontrivial in the sense it doesn’t border a disc in $M_{s}\cap B(p,r)$; if not below we can consider the next lowest disc and so on.
@@
Considering one of the boundary components $\gamma$ of $A_{1}$, we may clearly find an embedded disc $D$ in $N$ disjoint from $M_{s}$ with boundary $\gamma$. If $\gamma$ is nonseparating or splits $M_{s}$ into two components of genus greater than zero then $D$ will be a compressing disc because $\gamma$ will be homotopically nontrivial in $M$, which gives a contradiction. If the number of bridge points connected to the top sheet (i.e. where $B(p_{i},r_{i})$ are the bridge points and neighborhoods, the connected component of $M\setminus\cup_{i}B(p_{i},r_{i})$ containing $A_{1}$) is greater than one then $\gamma$ is nonseparating, so that we are done; let us suppose that $p$ is the only bridge point where the top sheet connects to the lower sheets then. If $\Sigma$ is not an embedded sphere or projective plane, then $\gamma$ would disconnect $M$ into two components of nonzero genus, completing the argument in this case. If $\Sigma$ is a sphere or projective plane, a bit more arguing is required. As before, first suppose that $\Sigma$ is a sphere. Then, because $N$ has trivial $\pi_{2}$, we may homotope $\Sigma$ and hence $M_{s}$ to lie in a small coordinate chart; of course, because ${\mathbb{R}}^{3}$ is contractible, $M_{s}$ in this case cannot be incompressible. If it is a projective plane again, the boundary of its tubular neighborhood is a sphere, and we may proceed as in the first case. ∎
@@
Leaving the world of stable minimal surfaces, we now consider the use of our first convergence statement, Theorem 1.1, to essentially replace the tightening process in minmax. Recalling that in the setting of Corollary 1.10 setting $N\simeq S^{3}$ we may consider the standard one parameter sweepout $\Sigma_{s}$ of it by 22–spheres, where there are two singular values of the sweepout that are single points corresponding to the north and south poles under a fixed choice of diffeomorphism from $N$ to $S^{3}$. Perhaps the most natural thing to do then is to consider the flow of the sweepout under the flow, by which we mean the family of the flows $(\Sigma_{s})_{t}$ of the individual leaves $\Sigma_{s}$. Recalling that the width of any sweepout of $S^{3}$ by spheres is positive by the isoperimetric inequality, we see if we could show that if the flow will remain a sweepout, then Theorem 1.1 says that the flow of at least one leaf will not go extinct, and because $S^{3}$ is simply connected and embedded hypersurfaces in such spaces are orientable Proposition 3.7 gives the limit will be an embedded minimal sphere, as usual potentially with multiplicity. There are some issues with the approach though, unfortunately.
@@
The biggest issue in general appears to be that the level set flow of some of the leaves may fatten, but in our particular setting, this may be ruled out as we discuss below. There is also an issue with preserving the typical regularity assumptions one makes in the definition of sweepout. Recall that a sweepout of $N$ (here, 11–parameter) is a family of surfaces $\Sigma_{s}$, $s\in[0,1]$, which for finitely many $s$ is singular in a finite set of points and smooth elsewhere and whose union is all of $N$. Now, we know that for a fixed leaf $\Sigma_{s}$ the flow $(\Sigma_{s})_{t}$ will be smooth for almost all times $t$, but because we wish to consider the flow of the family, parameterized over the whole unit interval (which is uncountable so Baire category theorem cannot apply), it doesn’t seem clear if we can show the flow of the whole family will be smooth for almost all or even many times. The current state of the art also doesn’t rule out the size of the singular set of a flow being greater than a finite union of points, either.
@@
With respect to the discussion above, one may be able to proceed by generalizing the results of minmax theory to more general sweepouts, including those "generated" by the flow, but instead, we will take a somewhat more hands-on approach. Denote by $E(s)$ the extinction time of the flow of $\Sigma_{s}$; if this never occurs, of course, we set $E(s)=\infty$ so it is a function into the extended reals. We next claim that $E(s)$ is sequentially continuous. To see this, we first note that because each of the $\Sigma_{s}$ is a sphere, their singularities under the flow must be mean convex and so each of their flows is nonfattening – implying, for instance, that the almost regular flows $(\Sigma_{s_{i}})_{t}$ are well defined. Fixing an $s_{0}\in[0,1]$, nonfattening also implies that for a sequence $s_{i}\to s_{0}$, after potentially passing to a subsequence by Brakke compactness, $\lim\limits_{s_{j}\to s_{0}}(\Sigma_{s_{i}})_{t}\to(\Sigma_{s_{0}})_{t}$ as Brakke flows because $\lim\limits_{s_{j}\to s_{0}}(\Sigma_{s_{i}})_{t}$ is a weak flow associated to $\Sigma_{s_{0}}$. With this in mind, suppose then that there was a sequence $s_{i}\to s_{0}$ and $\epsilon>0$ for which $E(s_{0})<A-\epsilon$, where $A=\lim\limits_{i\to\infty}E(s_{i})$.
@@
At time $T=E(s_{0})$, by Colding and Minicozzi [ColdingMinicozzi16_SingularSet] $(\Sigma_{s_{0}})_{T}$ is the union of a finite number of compact embedded Lipschitz curves corresponding to cylindrical singularities along with a countable collection of points corresponding to spherical singularities. This gives that for $i$ sufficiently large we make take the mass of $(\Sigma_{s_{i}})_{T}$ to be as small as we like, so in particular the clearing out lemma says these flows must go extinct before $t=A-\epsilon/2$, giving a contradiction to show that $E(s)$ is upper semicontinuous. Similarly, $E(s)$ is lower semicontinuous, giving our claim of sequential continuity.
@@
Because $E(s)$ is sequentially continuous, by the compactness of the unit interval, $\sup\limits_{s\in[0,1]}E(s)=T$ is obtained for some time $s_{0}$, recycling notation. Now, as above, we have $(\Sigma_{s_{0}})_{T}$ must consist of the union of a finite number of compact embedded Lipschitz curves corresponding to cylindrical singularities, along with a countable collection of points corresponding to spherical singularities. Picking some smooth time $t^{*}$ right before $T$, the resolution of the mean convex neighborhood conjecture gives that $(\Sigma_{s_{0}})_{t_{*}}$ is mean convex and after rescaling is locally modeled on bowl solitons, round cylinders, or round spheres.
@@
To see why this is useful, first note that $s_{0}\in(0,1)$: this is because for $s$ very near 0 or $1$ that $\Sigma_{s}$ are approximately small concentric round spheres contained in a coordinate chart so that $\lim\limits_{s\to 0,1}E(s)=0$; on the other hand for a fixed smooth $\Sigma_{s}$ its smooth flow exists for a short but strictly positive time. Secondly, the flows $(\Sigma_{s})_{t}$ are pairwise disjoint in $s$ because $\Sigma_{s}$ are and almost regular flows are weak set flows. In particular, for $s-s_{0}$ sufficiently small $s$ Brakke regularity says that $(\Sigma_{s})_{t^{*}}$ is a graph over $(\Sigma_{s_{0}})_{t_{*}}$ laying on one side or the other of it, and we can find $s$ so that the flow at time $t^{*}$ is on either side. Hence, we see that for $s$ either slightly greater or smaller than $s_{0}$, $(\Sigma_{s})_{T}$ must not be singular, or in other words, that their flow lasts at least slightly after time $T$, giving a contradiction to the definition of $T$. ∎